# An Evaluation of N-Gram Selection Strategies for Regular Expression Indexing in Contemporary Text Analysis Tasks [Experiment, Analysis & Benchmark]

### v1

@@ 4.2.2. Approximation Techniques
The budgeted maximum set cover problem optimization problem is NP-hard (Lund and Yannakakis, 1994; Khuller et al., 1999). Besides, the search space is the product of number of possible query-data pairs, $|Q|\cdot|D|$ and the number of possible n-grams, $|G|$. By Assumption Assumption BEST-4, we assume $G=G(Q)$. When the query size and the dataset size is large, the number of candidate n-grams can be prohibitively large to select the n-grams using brute-force method. BEST employs several techniques to get an approximate result.
@@ 4.2.3. Complexity Analysis
First, the workload reduction technique is applied, reducing the query set $Q$ to a reduced set $|Q^{\prime}|=\frac{|Q|}{t}$ for some $t\geq 1$ and the candidate n-gram set being $G^{\prime}$. It is reasonable to assume that $|G^{\prime}|\simeq\frac{|G|}{t}$. Let $\mu$ be the average number of n-grams in a query and $T$ be the maximum number of iterations of the k-median algorithm. Then, the workload reduction time is $O(|Q|\cdot|Q^{\prime}|\cdot\mu\cdot T)$. We use a suffix tree to efficiently enumerate all possible n-grams in the queries, which will take $O(|G^{\prime}|)$ time to build, enumerate, and insert them to a hash-map. A similar clustering is then carried out for creating sub-problems for parallelism. Empirically, according to the original BEST paper, dividing the problem to sub-problems reduces both time and space complexity by approximately $5$ even if we run it on a single thread. The use of the two adjacency lists Q-G-list and G-D-list rather than two matrices reduces the space overhead for utility calculation from $O(|G|\cdot(|D|+|Q|))$ to $O(\mu\cdot(|D|+|Q|))$. By Assumption Assumption BEST-4, $\mu\ll|G|$, and therefore, the memory overhead during computation is small. Note that we still need $O(|G|\cdot(|D|+|Q|))$ time to build both data structures. The two adjacency lists contributes to the majority of extra space overhead for BEST. In each iteration, BEST will examine all remaining candidate n-grams and their benefit considering the index $I$ built so far. For each pair $(q,d)\in(Q^{\prime}\times D)$, it checks if the pair covered by each candidate n-gram $g$. The time complexity of each iteration is $O(|G^{\prime}|\cdot|Q^{\prime}|\cdot|D|)$. The overall time complexity for BEST including workload reduction and clustering can be expressed as $O(|Q|\cdot|Q^{\prime}|\cdot\mu\cdot T+|G|\cdot(|D|+|Q|)+\frac{|I|\cdot|G^{\prime}|\cdot|Q^{\prime}|\cdot|D|}{5})$, which simplifies to $O(\frac{|I|}{5t^{2}}\cdot|G|\cdot|Q|\cdot|D|)$.
@@ 4.3.1. Selection Strategy
Definition 4.7 (LPMS).
Coverage The coverage of an n-gram $g$ is defined as the ratio of the support of $g$ in the dataset $D$ and the support of $g$ in the query set $Q$, normalized by the n-gram length.
$cv(g)=\frac{s_{D}(g)}{|g|\cdot s_{Q}(g)}$
@@
subject to $Ax\geq b$
$0\leq x_{g}\leq 1\quad\forall g\in G$
@@ 4.3.2. Complexity Analysis
To calculate the space overhead, let’s examine the sizes of each component. Both coverage and the output of the linear program are vectors of size $|G_{i}|$. As previously discussed, $A$ has a size of $|Q|\cdot|G_{i}|$ and $b$ has a size of $|Q|$. Summing these, the space overhead for LPMS n-gram selection is $O(|Q|\cdot|G_{i}|)$. During the algorithm runtime, coverage is calculated using $s_{D}(g)$ and $s_{Q}(g)$. By utilizing coverage for each n-gram rather than cover, LPMS reduces the time complexity for constructing the coverage vector to $O(|G_{i}|\cdot(|D|+|Q|))$ for each iteration. The linear program runs in polynomial time $O(|G_{i}|^{2.5})$ (Vaidya, 1989). In practice, the size of the index key typically does not exceed 10, a number also used as the upper bound for n-gram size in FREE. Therefore, we can consider the small number of iterations as a constant, making the overall computational complexity of LPMS $O(\sum_{i\in[10]}|G_{i}|^{2.5}+|G_{i}|\cdot(|D|+|Q|))$.
@@ 5. Experiment Setup
Due to the absence of source code from the papers, we implemented the three n-gram selection methods ourselves. For each method, we also developed their corresponding query parsers (if any) and index structures.
@@ 5.1. Benchmark Framework
The end-to-end process follows a seven-step sequence, as illustrated in the framework diagram Figure 2. First, the user specifies inputs (dataset, regex queries, method name, and configurations), which are then processed by the framework. The selected n-gram strategy processes the workload, identifying optimal n-grams for index construction. These n-grams are used to build an index (e.g., inverted index or B+-tree). The index search plan compiles by examining the regex patterns and extracting the n-grams in their literals to create plans for index lookup, filtering candidate data points. The regex engine then verifies these candidate data points, discarding false positives, and generating metrics to quantify performance. Finally, the results are aggregated and exported. The framework operates in a pipeline architecture divided into three phases: input processing, index construction, and regex evaluation. The components of each phase are grouped within dashed squares in Figure 2.
@@
Index Construction and Searching. In the index construction phase, one of the three n-gram selection strategies is applied to the workload. The selected n-grams are used to build an index, such as an inverted index or a B+-tree. This phase leverages multi-threaded execution if specified by the user. After the index is built, the index search plan is compiled if necessary. Each strategy has its own corresponding type of index structure and index search plan, as shown in the blue dashed bounding box in Figure 2. FREE and LPMS use inverted indexes, while BEST uses a B+-tree for the index structure. During index lookup, all n-grams in both the set of regex query literals and index keys are extracted. Only FREE requires an additional intermediate step to generate an index search tree.
@@
Table 2. Workload $W=(Q,D,\Sigma)$ statistics. We use $\bar{|d|}$ to denote average data record size (in bytes) and $\bar{TP}$ to denote the average number of actual matches of all queries. Synthetic has two query sets – the first set of 500 queries that is used to build the index, and a second set of 100 queries that is used to test the impact of queries that are not indexed.
Workload $|Q|$ $|D|$ ||$\sum$|| ∑ | $\bar{|d|}$ $\bar{TP}$ Dataset Size
@@ 5.2. Workloads
We use several real-world text datasets and queries with varying numbers of queries, data strings, alphabet sizes, average string lengths, and average matches per query for a comprehensive analysis. We summarize the workload characteristics in Section 5.1. Among the workloads, three (Webpages, DBLP, and Prosite) are those used by the original FREE, BEST, and LPMS papers to evaluate their methods. The other two workloads (US-Accand SQL-Srvr) are more recent and feature larger datasets.
@@
Synthetic. To test the robustness of the n-gram selection methods in handling unseen queries, we build off the synthetic dataset in the original LPMS (Tsang and Chawla, 2011) paper. The dataset consists of 5000 strings each constructed with alphabet ‘A-P’. The string size follows a geometric distribution with $p=1/32$. The index building query set is generated by randomly selecting 10% from the dataset, and a random slice is used to create a regex lit1regexlit2. lit1 has 1-5 characters, lit2 has 0-5 characters, and regex matches any $m$ characters where $1\leq m\leq 50$. During the query phase, we generate regexes of a similar format from 2% of the data records.
@@ 5.3. Metrics
Workload Matching Time ($T_{Q}$). We measure the time to run the workload, denoted by $T_{Q}$, to demonstrate the runtime gains provided by the index.
@@
Index Size ($S_{I}$). We measure the size of the index constructed, which is the total size of the n-gram keys, the posting lists, and necessary index data structure (such as the B+-tree).
@@ 5.4. Hardware and Implementation Details
We implemented FREE and its query parser, integrating it with an inverted index. For LPMS, we also used the same inverted index structure to ensure consistency. We implemented BEST with a B+-tree as the supporting index data structure. We used Gurobi Optimizer (Gurobi Optimization, LLC, 2024) as the linear program solver. All experiments were conducted on an Azure Standard_E32-16ds_v5 machine with an Intel(R) Xeon(R) Platinum 8370C CPU @ 2.80GHz, 16 vCPUs, 256 GB of memory, and 1 TB of disk storage. Our benchmarking framework is written in C++17 and compiled with the -O3 flag. Google’s RE2 (release version 2022-06-01) (Google, [n.d.]) as the regex engine for regex evaluation.
Each of the three methods has different configurable parameters. For BEST and FREE, we vary the selectivity threshold $c$ by picking values between 0.01 and 0.7. For FREE, we also vary the upper bound of n-gram length, $\max_{n}$, among values: 2, 4, 6, 8, and 10. Another parameter we vary in FREE is whether the selected n-gram set is prefix minimal or pre-suf minimal. We set a hard threshold of 3 hours for all experimental runs. If BEST fails to finish within this time limit, we reduce the workload by a factor of 0.1%, 0.2%, 0.5%, and 0.85%. LPMS allows us to configure whether to use deterministic relaxation (LPMS-D) or random relaxation (LPMS-R) of the linear programs. Similarly, we set a maximum number of n-grams for LPMS in case a workload takes too long to run. For all configurations of all methods, we use 16 threads. To fairly assess the methods, we compare the resulting indexes with a similar number of n-grams selected. Since FREE and LPMS do not natively support dynamic constraints on the index size, we implement an early-stopping mechanism for all three methods. By providing an optional parameter of max number of keys, the n-gram selection methods will stop once the limit is reached. We choose the values of max number of keys on a case-by-case basis for each workload, as specified in Section 6.1.
@@ 6. Experiments
In this section, we run the aforementioned workloads on the three indexes with varying parameters. In particular, we aim to answer the following questions:
@@
    How does each method perform on different datasets?
@@ 6.1. Query Performance
We determine the effectiveness of the index by evaluating the filtering effectiveness of the selected set of n-grams. The higher the precision of an index, the more effective the set of n-grams selected. Each configuration of a selection method generates a specific set of n-grams of varying sizes. To ensure a fair comparison, we compare selected sets of similar sizes, as indexing all possible n-grams would yield the highest precision for the workload. For each n-gram selection method under a specific key number constraint $K$ $s.t.$ $|I|\leq K$, we select the configuration with the highest precision. Note that throughout the section, we report the numbers for space usage and query performance for the configuration that achieves the best possible precision given a constraint on $K$.
@@ 6.1.1. Workload DBLP
Taking the DBLP workload as an example, we run BEST, FREE, and LPMS with varying parameters. After obtaining the results, we compare the performance of each n-gram selection method based on the upper bounds of the number of n-grams (i.e. index entries in $I$). Specifically, for this workload, we examine indices with a number of keys is at most $5000$. We summarize the results in Section 6.1.1, where the metric columns are variables introduced in Section 5.3.
Table 3. Index cost and query performance on DBLP.
$K$ Method $T_{Q}$ s $T_{I}$ s $S_{Q}$ $S_{I}$ $Prec$
150 BEST 19.7 533 2.194 33.433 0.235
FREE 25.3 2 0.171 37.299 0.051
LPMS 19.8 3 0.155 34.028 0.048
500 BEST 15.2 8762 2.670 54.306 0.242
FREE 20.6 2 0.242 66.774 0.165
LPMS 17.7 6 0.158 40.657 0.070
1000 BEST 15.2 8762 2.670 54.306 0.242
FREE 18.7 1 0.287 75.354 0.219
LPMS 17.0 808 1.309 41.499 0.071
2000 BEST 15.2 8762 2.670 54.306 0.242
FREE 17.4 1 0.264 68.442 0.321
LPMS 17.0 808 1.309 41.499 0.071
3000 BEST 15.2 8762 2.670 54.306 0.242
FREE 16.9 2 0.304 83.069 0.524
LPMS 17.0 808 1.309 41.499 0.071
4000 BEST 15.2 8762 2.670 54.306 0.242
FREE 16.9 2 0.304 83.069 0.524
LPMS 17.0 808 1.309 41.499 0.071
5000 BEST 15.2 8762 2.670 54.306 0.242
FREE 16.6 2 0.316 85.266 0.535
LPMS 17.0 808 1.309 41.499 0.071
[…]
When relaxing the index size constraint to 1000 n-grams, FREE ($\max_{n}=4$,$c=0.5$) generated an index of size 810 that achieved a precision of 0.219, slightly lower than BEST($c=0.5$). However, this configuration has a construction time of around 1 second, which is several orders of magnitude lower than both BEST and LPMS-D. FREE also consumes significantly less memory during computation. It uses 0.287 GB, whereas LPMS requires 1.3 GB and BEST 2.67 GB of memory to store the data and perform computation. For LPMS, with more n-grams selected, its second configuration achieves 10 times less index construction time and higher precision compared to using a smaller number of keys, resulting in a workload evaluation performance of 17.1 seconds, which is similar to the other methods. Although LPMS selects 929 keys to index, which is larger than the 810-key inverted index by FREE, it still has a smaller index size. This again demonstrates that LPMS tends to select n-grams with low frequencies in the dataset.
Figure 3. Index sizes of indices with different key numbers built by the three methods on DBLP workload.
As the number of n-grams allowed in the index increases, starting from 2000, more indices from different configurations of FREE fall into this range. These configurations have low selectivity thresholds ($c<0.15$) and $n=2$ compared to the best index generated by FREE when $K=1000$, selecting n-grams that do not cover enough strings. This is demonstrated by the smaller sizes of these indices in Figure 3 even as number of keys increase, representing their short posting lists. The best configurations of FREE($\max_{n}=2$,$c=0.2$) achieve higher precision than BEST. The precision increases from 0.321 at 2000 keys to 0.535 at 5000 keys upper limit by FREE($\max_{n}=4$,$c=0.1$), without a significant increase in construction overhead. As expected, the index size grows with the number of keys. However, the improvement in workload performance is less significant compared to the change in precision. In the DBLP workload, each string in the dataset is short, resulting in very quick regex evaluation.
Despite its precision, BEST achieves the best workload query performance. The index built by BEST has the smallest number of index keys, totaling 363. In comparison, a similar precision index by FREE requires 1440 keys. A larger index incurs greater index lookup overhead, which is significant when the dataset consists of short strings that match quickly after passing the filter. Another source of overhead for FREE comes from traversing its query plan trees, which are optimized for regexes with alternation (logical OR).
@@
FREE performs best on the DBLP workload. Generally, if the workload has a large number of queries that are not skewed (i.e., all very similar and covering a small subset of the dataset) and the strings in the dataset are not long, it is best to choose FREE. When the query set is large, BEST and LPMS require significant computation time and memory, owing to their higher time and space complexities. Additionally, for a large-sized and balanced query set, frequent n-grams in the datasets are likely to be covered in the queries as well. Therefore, FREE benefits from these two aspects and outperforms the other two methods.

### v2

@@ 4.2.2. Approximation Techniques
The budgeted maximum set cover problem optimization problem is NP-hard (Lund and Yannakakis, 1994; Khuller et al., 1999). Besides, the search space is the product of number of possible query-data pairs, $|Q|\cdot|D|$ and the number of possible n-grams, $|G|$. When the query size and the dataset size is large, the number of candidate query-data pairs can be prohibitively large to select the n-grams using brute-force method. BEST employs several techniques to get an approximate result.
@@ 4.2.3. Complexity Analysis
First, the workload reduction step shrinks the query set from $|Q|$ to $|Q^{\prime}|=\frac{|Q|}{t}$, with the candidate n-gram set $G^{\prime}$ also reduced proportionally. Assuming average query length $\mu$ and $T$ iterations of k-median, this step takes $O(|Q|\cdot|Q^{\prime}|\cdot\mu\cdot T)$ time. A suffix tree is used to enumerate all n-grams in $O(|G^{\prime}|)$ time. Next, clustering enables parallelism and, per the original BEST paper, reduces time and space by $\sim$5 $\times$ even on a single thread. Using adjacency lists (Q-G-list and G-D-list) instead of matrices cuts space from $O(|G|\cdot(|D|+|Q|))$ to $O(\mu\cdot(|D|+|Q|))$, which is small since $\mu\ll|G|$. However, building these lists still takes $O(|G|\cdot(|D|+|Q|))$ time. By Assumption Assumption BEST-4, $\mu\ll|G|$, and therefore, the memory overhead during computation is small. Note that we still need $O(|G|\cdot(|D|+|Q|))$ time to build both data structures. The two adjacency lists contributes to the majority of extra space overhead for BEST. In each iteration, BEST will examine all remaining candidate n-grams and their benefit considering the index $I$ built so far. For each pair $(q,d)\in(Q^{\prime}\times D)$, it checks if the pair covered by each candidate n-gram $g$. In each iteration, BEST evaluates all remaining n-grams against $(q,d)\in Q^{\prime}\times D$, costing $O(|I|\cdot|G(Q)|\cdot|Q|\cdot|D|)$.
@@ 4.3.1. Selection Strategy
Definition 4.7 (LPMS Coverage).
The coverage of an n-gram $g$ is defined as the ratio of the support of $g$ in the dataset $D$ and the support of $g$ in the query set $Q$, normalized by the n-gram length: $cv(g)={s_{D}(g)}/{|g|\cdot s_{Q}(g)}$
@@
subject to $Ax\geq b;0\leq x_{g}\leq 1\quad\forall g\in G$
@@ 4.3.2. Complexity Analysis
To calculate the space overhead, let’s examine the sizes of each component. Both coverage and the output of the linear program are vectors of size $|G_{i}|$. As previously discussed, $A$ has a size of $|Q|\cdot|G_{i}|$ and $b$ has a size of $|Q|$. Summing these, the space overhead for LPMS n-gram selection is $O(|Q|\cdot|G_{i}|)$. During the algorithm runtime, coverage is calculated using $s_{D}(g)$ and $s_{Q}(g)$. By utilizing coverage for each n-gram rather than cover, LPMS reduces the time complexity for constructing the coverage vector to $O(|G_{i}|\cdot(|D|+|Q|))$ for each iteration. The linear program runs in polynomial time $O(|G_{i}|^{2.5})$ (Vaidya, 1989). In practice, the size of the index key typically does not exceed 10, a number also used as the upper bound for n-gram size in FREE. Therefore, we can consider the small number of iterations as a constant, making the overall computational complexity of LPMS $O(\sum_{i\in[10]}\{|G_{i}|^{2.5}+|G_{i}|\cdot(|D|+|Q|)\})$.
4.4. VGGraph
The n-gram selection strategy from VGGraph (Qiu et al., 2025) combines the insights from FREE and BEST to select the n-grams.
4.4.1. Selection Strategy
The idea is to select the most cost-effective n-gram from a set $S$ to cover characters in an uncovered set $E$. Each n-gram is evaluated based on a cost-efficiency ratio (a notion similar to cover from Definition 4.3), where the cost is size of the inverted index for the n-gram and the efficiency is the number of characters it helps cover in $E$. VGGraph approximate a perfect set cover with efficiency by adopting FREE’s idea of Usefulness, and only extend the useful n-grams. By maintaining the uncovered set $E$ that is continuously updated as the n-grams are picked, VGGraph ensures that prefix-free n-grams are selected preferentially.
4.4.2. Complexity Analysis
VGGraph’s index key selection consists of two steps: (1) recursive extension of grams exceeding the frequency threshold $c$, and (2) greedy n-gram selection via the set-cover approximation. Given a dataset $D$, the complexity of recursively extending grams is $O(|D|\cdot q_{\min})$, as each character position is indexed exactly once and each gram is extended at most once. The greedy set-cover step selects, for each regex, a minimal subset of grams from the candidate set $G(Q)$. For a regex with literal size $n$, the greedy selection has a complexity of $O(|G(Q)|\cdot n)$ per query. The total runtime complexity for key selection, combining both steps, is thus $O(|D|\cdot q_{\min}+|Q|\cdot|G(Q)|\cdot n)$. The corresponding runtime space complexity during key selection is $O(|D|+|G(Q)|)$, dominated by storing candidate grams and their counts.
@@ 5. Experiment Setup
Due to the absence of source code from the papers, we implemented the three n-gram selection methods ourselves. For a fair comparison of the n-gram selection methods, we use the same index structure (inverted index) and same basic query matcher for all experiments. We include the implementations of other index structures and query plans per the original papers in our codebase for reference.
@@ 5.1. Benchmark Framework
The end-to-end process follows a seven-step sequence, as illustrated in the framework diagram Figure 1. t begins with user-specified inputs (dataset, regex queries, method, and configurations), which are processed by the framework. The selected n-gram strategy analyzes the workload to identify optimal n-grams for index construction. These n-grams are then used to build an index (e.g., inverted index or B+-tree). Next, the index search plan extracts n-grams from regex literals to guide candidate filtering. The regex engine verifies these candidates, discards false positives, and computes performance metrics. Finally, results are aggregated and exported. The pipeline is organized into three phases—input processing, index construction, and regex evaluation—with components grouped accordingly in Figure 1.
@@
Index Construction and Searching. In the index construction phase, one of the three n-gram selection strategies is applied to the workload. The selected n-grams are used to build an index. This phase leverages multi-threaded execution if specified by the user. After the index is built, the index search plan is compiled if necessary. An inverted index with data string IDs is built with the selected n-grams. In the simple query plan, posting lists corresponding to the set of n-grams that exists in the regex query literals are unioned together later for full regex evaluation.
@@
Table 2. Workload $W=(Q,D,\Sigma)$ statistics. We use $\bar{|d|}$ to denote average data record size (in bytes) and $\bar{TP}$ to denote the average number of actual matches of all queries. Robust has two query sets – the first set of 500 queries that is used to build the index, and a second set of 100 queries that is used to test the impact of queries that are not indexed.
Workload $|Q|$ $|D|$ $|\sum|$ $\bar{|d|}$ $\bar{TP}$ Dataset Size
@@ 5.2. Workloads
We use several real-world text datasets and queries with varying numbers of queries, data strings, alphabet sizes, average string lengths, and average matches per query for a comprehensive analysis. We summarize the workload characteristics in Section 5.1. Among the workloads, three (Webpages, DBLP, and Prosite) are those used by the original FREE, BEST, and LPMS papers to evaluate their methods. The other four workloads (US-Acc, SQL-Srvr, Enron, ADX) are more recent and feature larger datasets.
@@
Enron. This is a production workload consists of more than 50 thousand of real emails and 107 regex queries. The dataset is the May 7, 2015 Version of Enron dataset (Cohen, 2015), and the queries are from Avatar search engine (Jayram et al., 2006) collected by Chen et al. (Chen, 2012).
ADX. This is a real production workload consists of 890 millions of logs from Azure Data Explorer (ADX) and 18 regexes for an log analysis task on this dataset by an data scientist.
Synthetic. This workload is based on the method described in the paper proposing LPMS (Tsang and Chawla, 2011). The workload consists of 50,000 data records with a fixed alphabet size of 16. We varied the query set size across $[500,1000,2000,5000]$ and explored different levels of query selectivity ranging from 0.001 to 0.5.
@@ 5.3. Metrics
Query Time ($T_{Q}$). We measure the time to run the workload, denoted by $T_{Q}$, to demonstrate the runtime gains provided by the index.
@@
Index Size ($S_{I}$). We measure the size of the index constructed, which is the total size of the n-gram keys, the posting lists, and necessary index data structure (such as the inverted index).
@@ 5.4. Hardware and Implementation Details
We implemented n-gram selection methods FREE, BEST, LPMS, VGGraph, and Trigrams with the index construction and matching methods mentioned in the benchmarking framework. We implemented FREE’ query parser and BEST with a B+-tree in the codebase as well. We used Gurobi Optimizer (Gurobi Optimization, LLC, 2024) as the linear program solver. All experiments were conducted on an Azure Standard_E32-16ds_v5 machine with an Intel(R) Xeon(R) Platinum 8370C CPU @ 2.80GHz, 16 vCPUs, 256 GB of memory, and 1 TB of disk storage. Our benchmarking framework is written in C++17 and compiled with the -O3 flag. Google’s RE2 (release version 2022-06-01) (Google, [n.d.]) as the regex engine for regex evaluation.
Each method has its own configurable parameters. For BEST and FREE, we vary the selectivity threshold $c$ between 0.01 and 0.7. FREE also varies the maximum n-gram length ($\max_{n}\in\{2,4,6,8,10\}$) and whether the selected n-grams are prefix-minimal or pre-suf-minimal. All experiments are capped at 3 hours; if BEST exceeds this, we reduce the workload to 0.1–0.85%. LPMS supports both deterministic (LPMS-D) and randomized (LPMS-R) relaxations. We also cap the number of n-grams for LPMS if runtime is excessive. All methods run with 16 threads. To ensure fair comparison, we match index sizes across methods. Since FREE and LPMS don’t natively support key limits, we implement early stopping via a max-key parameter for all three methods. By providing an optional parameter of max number of keys, the n-gram selection methods will stop once the limit is reached. Each configuration of a selection method generates a specific set of n-grams of varying sizes. To ensure a fair comparison, we compare selected sets of similar sizes, as indexing all possible n-grams would yield the highest precision for the workload. Since each methods has different configurations, and those configuration values interfere with each other in a non-linear way, isolating one configuration and study its behavior as value changes is not meaningful. For each n-gram selection method under a specific key number constraint $K$ $s.t.$ $|I|\leq K$, we select the configuration with the highest precision. We choose the values of max number of keys, $K$, on a case-by-case basis for each workload, as specified in Section 6.
@@ 6. Experiments
In this section, we run the aforementioned workloads on the indexes build from different n-gram selection methods with varying parameters. In particular, we aim to answer the following questions:
@@
    What is the usefulness of each n-gram selection method? We evaluate this through precision and $T_{Q}$.
@@
    What is the index construction overhead for each method? We evaluate this through $T_{I}$, $S_{Q}$, and $S_{I}$.
  3. Q.3
@@
We answer question Q.1 in Section 6.1 and question Q.2 in Section 6.2. To answer question Q.3, we discuss the impact of characteristics of different real-world workloads on index construction overhead in Section 6.2 and on index effectiveness in Section 6.1. We also use present a more thorough ablation study on impact of workload parameters such as the query set size $|Q|$, dataset size $|D|$, , query selectivity, etc. in Section 6.3.
(a) Precision
(b) Query Time $T_{Q}$ (s)
Figure 2. DBLP.
@@ 6.1. Query Performance
To address question Q.1, we compare the performance gains from using indices built with n-grams selected by the three methods, focusing on filtering precision. Higher precision indicates a more effective n-gram set. We also measure the regex query runtime $T_{Q}$, which reflects precision since the index structure and matching strategy remain constant.
Generally, indexing more n-grams improves precision. However, each method behaves differently. BEST often starts with the highest precision when the key budget $K$ is small, but its precision gains plateau due to time constraints (e.g., 3-hour index build). In contrast, FREE begins with lower precision—especially for small query sets—but improves significantly as $K$ increases, often achieving the highest precision overall. Precision of VGGraph is low when number of keys is small, but quickly rises as number of keys gets larger. LPMS has similar precision as BEST when the numbers of key is small and usually grows insignificantly as the number of n-grams selected increases. Trigrams typically has very low precision when the number of index keys fixed is low. We now investigate the methods’ precision for different workloads with varying characteristics.
@@ 6.1.1. Workload DBLP
We evaluate BEST, FREE, LPMS, and VGGraph across various input configurations, and applying the same key limit to Trigrams. Performance is compared based on index size, focusing on configurations with at most 5000 n-grams.
As shown in Figure 2(a), BEST($c=0.5$) achieves a precision of 0.235 for $K\leq 150$, outperforming FREE($\max_{n}=2$, $c=0.5$) and LPMS-D, which yield much lower precision (0.051 and 0.048, respectively) at a similar index size. At $K\leq 500$, the 363 n-grams selected by BEST($c=0.5$) achieves significantly better precision compared to LPMS-D. With similar index construction overhead, FREE($\max_{n}=2$,$c=0.5$) achieves higher precision than LPMS-D. LPMS tends to select infrequent n-grams, which may not benefit overall workload filtering. Although BEST improves slightly with more keys (e.g., +0.007 precision and +4.5s query performance from $K=150$ to $K=500$), gains are modest. Meanwhile, VGGraph($\max_{n}=2$,$c=0.5$) builds an index with 215 n-grams but with a low precision of 0.035.
When relaxing the index size constraint to 1000 n-grams, FREE ($\max_{n}=4$,$c=0.5$) generated an index of size 810 that achieved a precision of 0.219, slightly lower than BEST($c=0.5$). VGGraph($max_{n}=3$, $c=0.5$)’s precision increase is insignificant. As the number of n-grams allowed in the index increases, starting from 2000, more indices from different configurations of FREE fall into this range. The best configurations of FREE($\max_{n}=2$,$c=0.2$) achieve higher precision than BEST. The precision increases from 0.321 at 2000 keys to 0.524 at 4000 keys upper limit by FREE($\max_{n}=4$,$c=0.15$). Precision of VGGraph increases significantly from 0.167 to 0.725 as number of keys increases from 1014 to 3681.
DBLP runs much faster with all index configurations than the baseline (i.e. using no indexes at all) of 99 seconds. Even when Trigrams indexes are built using only the first few trigrams—yielding low precision between 0.006 and 0.017, they still achieves performance improvement of $1.8\times$ to $3.6\times$. Looking at Figure 2(b), however, the improvement in workload performance is significant compared to the change in precision. The precision of FREE increases around 10 times as $K$ increases from 150 to 3000; its query time decreases by 41%. When $K=150$, the precision of BEST is significantly larger than FREE and LPMS in Figure 2(a), the performance difference for BEST and LPMS is not significantly different. In the DBLP workload, each string in the dataset is short, resulting in very quick regex evaluation, and therefore the query time reduction brought by precision increase is less significant. Trigrams’s precision increases as number of keys increases. When indexing all possible 28,275 trigrams, we can get a high precision of 0.97. However, the query time is 14.7s, similar to BEST with less than 500 keys. This suggests that indexing all trigrams is not cost-efficient in terms of storage.
@@
FREE performs best on the DBLP workload. Generally, if the workload has a large number of queries that are not skewed (i.e., all very similar and covering a small subset of the dataset) and the strings in the dataset are not long, it is best to choose FREE. VGGraph significantly improve its precision only when the number of keys increase. When the query set is large, BEST and LPMS require significant computation time and memory, owing to their higher time and space complexities. Additionally, for a large-sized and balanced query set, frequent n-grams in the datasets are likely to be covered in the queries as well.
Table 3. Index cost and query performance on Webpages.
$K$ Method $T_{Q}$ s $T_{I}$ s $S_{Q}$ GB $S_{I}$ MB PrecPrec
[…]